\newcommand{\etalchar}[1]{$^{#1}$}
\begin{thebibliography}{CKM{\etalchar{+}}19}

\bibitem[BBDS12]{BlockiBDS12}
Jeremiah Blocki, Avrim Blum, Anupam Datta, and Or~Sheffet.
\newblock The johnson-lindenstrauss transform itself preserves differential
  privacy.
\newblock In {\em 53rd IEEE Symposium on Foundations of Computer Science},
  FOCS'12, pages 410--419, New Brunswick, NJ, USA, 2012.

\bibitem[BBDS13]{BlockiBDS13}
Jeremiah Blocki, Avrim Blum, Anupam Datta, and Or~Sheffet.
\newblock Differentially private data analysis of social networks via
  restricted sensitivity.
\newblock In {\em 4th ACM Conference on Innovations in Theoretical Computer
  Science}, ITCS '13, pages 87--96, Berkeley, CA, USA, 2013. ACM.

\bibitem[BCSZ18]{BorgsCSZ18}
Christian Borgs, Jennifer~T. Chayes, Adam~D. Smith, and Ilias Zadik.
\newblock Revealing network structure, confidentially: Improved rates for
  node-private graphon estimation.
\newblock In {\em 59th Annual IEEE Symposium on Foundations of Computer
  Science}, FOCS '18, pages 533--543, Paris, France, 2018.

\bibitem[BS19]{BunS19}
Mark Bun and Thomas Steinke.
\newblock Smooth sensitivity, revisited.
\newblock Manuscript, 2019.

\bibitem[CD18]{CummingsD18}
Rachel Cummings and David Durfee.
\newblock Individual sensitivity preprocessing for data privacy.
\newblock {\em arXiv preprint arXiv:1804.08645}, 2018.

\bibitem[CKM{\etalchar{+}}19]{CanonneKMUZ19}
Cl\'{e}ment~L. Canonne, Gautam Kamath, Audra McMillan, Jonathan Ullman, and
  Lydia Zakynthinou.
\newblock Differentially private identity testing for multivariate
  distributions.
\newblock Manuscript, 2019.

\bibitem[DMNS06]{DworkMNS06}
Cynthia Dwork, Frank McSherry, Kobbi Nissim, and Adam Smith.
\newblock Calibrating noise to sensitivity in private data analysis.
\newblock In {\em Proceedings of the 3rd Conference on Theory of Cryptography},
  TCC '06, pages 265--284, Berlin, Heidelberg, 2006. Springer.

\bibitem[GRU12]{GuptaRU12}
Anupam Gupta, Aaron Roth, and Jonathan Ullman.
\newblock Iterative constructions and private data release.
\newblock In {\em 9th IACR Theory of Cryptography Conference}, TCC '12, pages
  339--356, Taormina, Italy, 2012. Springer.

\bibitem[HLMJ09]{HayLMJ09}
Michael Hay, Chao Li, Gerome Mikalu, and David~D. Jensen.
\newblock Accurate estimation of the degree distribution of private networks.
\newblock In {\em Proceedings of the 9th IEEE International Confernece on Data
  Mining}, ICDM'09, pages 169--178, Miami, FL, USA, 2009.

\bibitem[KNRS13]{KasiviswanathanNRS13}
Shiva~P. Kasiviswanathan, Kobbi Nissim, Sofya Raskhodnikova, and Adam~D. Smith.
\newblock Analyzing graphs with node differential privacy.
\newblock In {\em 10th IACR Theory of Cryptography Conference}, TCC '13, pages
  457--476, Tokyo, Japan, 2013. Springer.

\bibitem[KRSY14]{KarwaRSY14}
Vishesh Karwa, Sofya Raskhodnikova, Adam~D. Smith, and Grigory Yaroslavtsev.
\newblock Private analysis of graph structure.
\newblock {\em ACM Transactions on Database Systems}, 39(3):22:1--22:33, 2014.

\bibitem[KS16]{KarwaS16}
Vishesh Karwa and Aleksandra Slavkovi{\'c}.
\newblock Inference using noisy degrees: Differentially private $\beta$-model
  and synthetic graphs.
\newblock {\em Annals of Statistics}, 44(1):87--112, 2016.

\bibitem[NRS07]{NissimRS07}
Kobbi Nissim, Sofya Raskhodnikova, and Adam Smith.
\newblock Smooth sensitivity and sampling in private data analysis.
\newblock In {\em Proceedings of the 30th annual ACM Symposium on Theory of
  Computing, STOC}, pages 75--84, 2007.

\bibitem[RS16]{RaskhodnikovaS16}
Sofya Raskhodnikova and Adam~D. Smith.
\newblock Lipschitz extensions for node-private graph statistics and the
  generalized exponential mechanism.
\newblock In {\em 57th Annual IEEE Symposium on Foundations of Computer
  Science}, FOCS '16, pages 495--504, New Brunswick, NJ, USA, 2016.

\bibitem[XCT14]{XiaoCT14}
Qian Xiao, Rui Chen, and Kian-Lee Tan.
\newblock Differentially private network data release via structural inference.
\newblock In {\em 20th ACM International Conference on Knowledge Discovery and
  Data Mining}, KDD'14, pages 911--920, 2014.

\end{thebibliography}

\def\comments{0}

\documentclass[11pt]{article}
\usepackage{preamble}
\usepackage{lipsum}

\newcommand{\wt}{\mathsf{wt}}
\newcommand{\val}{\mathsf{val}}
\renewcommand{\epsilon}{\varepsilon}
\newcommand{\ER}{Erd\H{o}s-R\'{e}nyi}
\newcommand{\ad}{\bar{d}}
\newcommand{\kstar}{k^*}
\newcommand{\kprime}{k'}
\newcommand{\kprimeprime}{k''}

\newcommand{\cN}{\mathcal{N}}

\newcommand{\Far}{\mathsf{Far}}

\newcommand{\mypar}[1]{\medskip\noindent\textbf{#1}}

\title{Efficiently Estimating Erd\H{o}s-R\'{e}nyi Graphs with \\ Node Differential Privacy}
\author[1]{Adam Sealfon}
\author[2]{Jonathan Ullman}
\affil[1]{EECS and CSAIL, MIT}
\affil[2]{Khoury College of Computer Sciences, Northeastern University}

\begin{document}

\maketitle


\begin{abstract}
We give a simple, computationally efficient, and node-differentially-private algorithm for estimating the parameter of an \ER~graph---that is, estimating $p$ in a $G(n,p)$---with near-optimal accuracy.  Our algorithm nearly matches the information-theoretically optimal exponential-time algorithm for the same problem due to Borgs et al.~(FOCS 2018).  More generally, we give an optimal, computationally efficient, private algorithm for estimating the edge-density of any graph whose degree distribution is concentrated on a small interval. 
\end{abstract}

\section{Introduction}

Network data modeling individuals and relationships between individuals are increasingly central in data science.  However, while there is a highly successful literature on differentially private statistical estimation for traditional iid data, the literature on estimating network models is far less well developed.

Early work on private network data focused on \emph{edge-differential-privacy}, in which the algorithm is required to ``hide'' the presence or absence of a single edge in the graph (see, e.g.~\cite{NissimRS07,HayLMJ09,KarwaRSY14,GuptaRU12,BlockiBDS12,XiaoCT14,KarwaS16}, and many others).  A more desirable notion of privacy is \emph{node-differential privacy (node-DP)}, which requires the algorithm to hide the presence or absence of an arbitrary set of edges incident on a single node.  Although node-DP is difficult to achieve without compromising accuracy, the beautiful works of Blocki \etal~\cite{BlockiBDS13} and Kasiviswanathan \etal~\cite{KasiviswanathanNRS13} showed how to design accurate node-DP estimators for many interesting graph statistics via \emph{Lipschitz extensions.}  However, many of the known constructions of Lipschitz extensions require exponential running time, and constructions of computationally efficient Lipschitz extensions~\cite{RaskhodnikovaS16,CummingsD18,CanonneKMUZ19} lag behind.  As a result, even for estimating very simple graph models, there are large gaps in accuracy between the best known computationally efficient algorithms and the information theoretically optimal algorithms.

In this work we focus on what is arguably the simplest model of network data, the \emph{\ER~graph}.  In this model, denoted $G(n,p)$, we are given a number of nodes $n$ and a parameter $p \in [0,1]$, and we sample an $n$-node graph $G$ by independently including each edge $(i,j)$ for $1 \leq i < j \leq n$ with probability $p$.  The goal is to design a node-DP algorithm that takes as input a graph $G \sim G(n,p)$ and outputs an estimate $\hat{p}$ of the edge density parameter $p$.

Surprisingly, until an elegant recent work of Borgs \etal~\cite{BorgsCSZ18}, the optimal accuracy for estimating the parameter $p$ in a $G(n,p)$ via node-DP algorithms was unknown.  Although that work essentially resolved the optimal accuracy of node-DP algorithms, their construction is again based on generic Lipschitz extensions, and thus results in an exponential-time algorithm, and, in our opinion, gives little insight for how to construct an efficient estimator with similar accuracy.

The main contribution of this work is to give a simple, polynomial-time estimator for \ER~graphs whose error very nearly matches that of Borgs \etal's estimator, and indeed matches it in a wide range of parameters.  We achieve this by giving a more general result, showing how to optimally estimate the edge-density of any graph whose degree distribution is concentrated in a small interval.

\subsection{Background: Node-Private Algorithms for \ER~Graphs}

Without privacy, the optimal estimator is simply to output the \emph{edge-density} $p_{G} = |E|/\binom{n}{2}$ of the realized graph $G \sim G(n,p)$, which guarantees
$$
\ex{G}{(p - p_{G})^2} = \frac{p(1-p)}{\binom{n}{2}}.
$$
The simplest way to achieve $\eps$-node-DP is to add zero-mean noise to the edge-density with standard-deviation calibrated to its \emph{global-sensitivity}, which is the amount that changing the neighborhood of a single node in a graph can change its edge-density.  The global sensitivity of $p_{G}$ is $\Theta(1/n)$, and thus the resulting private algorithm $\cA_{\textit{na\"{i}ve}}$ satisfies
$$
\ex{G}{(p - \cA_{\textit{na\"{i}ve}}(G))^2} = \Theta\left( \frac{1}{\eps^2 n^2} \right)
$$
Note that this error is \emph{at least} on the same order as the non-private error, and can asymptotically dominate the non-private error.

Borgs \etal~\cite{BorgsCSZ18} gave an improved $\eps$-node-DP algorithm such that, when both $p$ and $\eps$ are $\gtrsim \log(n)/n$,
$$
\ex{}{(p - \cA_{\textit{bcsz}}(G))^2} = \underbrace{\frac{p(1-p)}{\binom{n}{2}}}_{\textrm{non-private error}} +~\underbrace{\tilde{O}\left( \frac{p}{\eps^2 n^{3}} \right)}_{\textrm{overhead due to privacy}}
$$
What is remarkable about their algorithm is that, unless $\eps$ is quite small (roughly $\eps \lesssim n^{-1/2}$), the first term dominates the error, in which case privacy comes essentially \emph{for free}.  That is, the error of the private algorithm is only larger than that of the optimal non-private algorithm by a $1+o(1)$ factor.  However, as we discussed above, this algorithm is not computationally efficient.

The only computationally efficient node-DP algorithms for computing the edge-density apply to graphs with small maximum degree~\cite{BlockiBDS13,KasiviswanathanNRS13,RaskhodnikovaS16}, and thus do not give optimal estimators for \ER~graphs unless $p$ is very small.

\subsection{Our Results}

Our main result is a computationally efficient estimator for \ER~graphs.

\jnote{Need to be careful about what the algorithm depends on.  I think either it needs to be given some bound on $p$ or else it will be a little weaker.}

\begin{thm}[\ER~Graphs, Informal] \label{thm:main-er}
There is an $O(n^2)$-time $\eps$-node-DP algorithm $\cA$ such that for every $n$ and every $p \gtrsim 1/n$ if $G \sim G(n,p)$ then
$$
\ex{G,A}{(p - \cA(G))^2} = \underbrace{\frac{p(1-p)}{\binom{n}{2}}}_{\textrm{non-private error}} +~~\underbrace{\tilde{O}\left( \frac{p}{\eps^2 n^{3}} + \frac{1}{\eps^4 n^4} \right)}_{\textrm{overhead due to privacy}}
$$
\end{thm}
The error of Theorem~\ref{thm:main-er} matches that of the exponential-time estimator of Borgs \etal~\cite{BorgsCSZ18} up to the additive $\tilde{O}(1/\eps^4 n^4)$ term, which is often not the dominant term in the overall error.  In particular, the error of our estimator is still within a 1+o(1) factor of the optimal non-private error unless $\eps$ or $p$ is quite small---for example, when $p$ is a constant and $\eps \gtrsim n^{-1/2}$.  

\medskip
Our estimator actually approximates the edge density for a significantly more general class of graphs than merely \ER~graphs.  Specifically, Theorem~\ref{thm:main-er} follows from a more general result for the family of \emph{concentrated-degree graphs}.  For $k \in \N$, define $\cG_{n,k}$ to be the set of $n$-node graphs such that the degree of every node is between $\ad - k$ and $\ad + k$, where $\ad = 2 |E| / n$ is the average degree of the graph.

\begin{thm}[Concentrated-Degree Graphs, Informal] \label{thm:main-cd}
For every $k \in \N$, there is an $O(n^2)$-time $\eps$-node-DP algorithm $\cA$ such that for every $n$ and every $G \in \cG_{n,k}$,
$$
\ex{\cA}{\left(p_{G} - \cA(G)\right)^2} = O\left( \frac{k^2}{\eps^2 n^4} + \frac{1}{\eps^4 n^4} \right)
$$
where $p_{G} = |E|/\binom{n}{2}$ is the empirical edge density of $G$.  
\end{thm}

Theorem~\ref{thm:main-er} follows from Theorem~\ref{thm:main-cd} by using the fact that for an \ER~graph, with overwhelming probability the degree of every node lies in an interval of width $\tilde{O}(\sqrt{pn})$ around the average degree.

The main technical ingredient in Theorem~\ref{thm:main-cd} is to construct a \emph{low sensitivity} estimator $f(G)$ for the number of edges.  The first property we need is that when $G$ satisfies the concentrated degrees property, $f(G)$ equals the number of edges in $G$.  The second property of the estimator we construct is that its \emph{smooth sensitivity}~\cite{NissimRS07} is low on these graphs $G$.  At a high level, the smooth sensitivity of $f$ at a graph $G$ is the most that changing the neighborhood of a small number of nodes in $G$ can change the value of $f(G)$.  Once we have this property, it is sufficient to add noise to $f(G)$ calibrated to its smooth sensitivity.  We construct $f$  by carefully reweighting edges that are incident on nodes that do not satisfy the concentrated-degree condition.

\medskip
Finally, we are able to show that Theorem~\ref{thm:main-cd} is optimal for concentrated-degree graphs.  In additional to being a natural class of graphs in its own right, this lower bound demonstrates that in order to improve Theorem~\ref{thm:main-er} we will need techniques that are more specialized to \ER~graphs.

\begin{thm}[Lower Bound, Informal] \label{thm:main-lb}
For every $n$ and $k$, and every $\eps$-node-DP algorithm $A$, there is some $G \in \cG_{n,k}$ such that
$$
\ex{A}{\left(p_{G} - \cA(G)\right)^2} = \Omega\left( \frac{k^2}{\eps^2 n^4} + \frac{1}{\eps^4 n^4} \right)
$$
The same bound applies to $(\eps,\delta)$-node-DP algorithms with sufficiently small $\delta \lesssim \eps$.
\end{thm}

\section{Preliminaries}

Let $\cG_{n}$ be the set of $n$-node graphs.  We say that two graphs $G,G' \in \cG_{n}$ are \emph{node-adjacent}, denoted $G \sim G'$, if $G'$ can be obtained by $G$ modifying the neighborhood of a single node $i$. That is, there exists a single node $i$ such that for every edge $e$ in the symmetric difference of $G$ and $G'$, $e$ is incident on $i$.  As is standard in the literature on differential privacy, we treat $n$ as a fixed quantity and define adjacency only for graphs with the same number of nodes.  We could easily extend our definition of adjacency to include adding or deleting a single node itself. 

\begin{definition}[Differential Privacy~\cite{DworkMNS06}]
A randomized algorithm $\cA \from \cG_{n} \to \cR$ is $(\eps,\delta)$-node-differentially private if for every $G \sim G' \in \cG_{n}$ and every $R \subseteq \cR$,
$$
\pr{}{\cA(G) \in R} \leq e^{\eps} \cdot \pr{}{\cA(G') \in R} + \delta
$$
If $\delta = 0$ we will simply say that $\cA$ is $\eps$-node-differentially private.  As we only consider node differential privacy in this work, we will frequently simply say that $\cA$ satisfies differential privacy.
\end{definition}

The next lemma is the basic composition property of differential privacy.
\begin{lem}[Composition~\cite{DworkMNS06}] \label{lem:dpcomp} If $\cA_1, \cA_2 \from \cG_{n} \to \cR$ are each $(\eps,\delta)$-node-differentially private algorithms, then the mechanism $\cA(G) = (\cA_1(G), \cA_2(G))$ satisfies $(2 \eps,2\delta)$-node-differential privacy.  The same holds if $\cA_2$ may depend on the output of $\cA_1$.
\end{lem}

We say that two graphs $G,G'$ are at \emph{node distance $c$} if there exists a sequence of graphs
$$
G = G_0 \sim G_1 \dots G_{c-1} \sim \dots G_{c} = G'
$$
The standard \emph{group privacy} property of differential privacy yields the following guarantees for graphs at node distance $c > 1$.
\begin{lem}[Group Privacy~\cite{DworkMNS06}] \label{lem:groupdp} If $\cA \from \cG_n \to \cR$ is $(\eps,\delta)$-node-differentially-private and $G,G'$ are at node-distance $c$ then for every $R \subseteq \cR$,
$$
\pr{}{\cA(G) \in R} \leq e^{c \eps}  \pr{}{\cA(G') \in R} + c e^{c \eps} \delta
$$
\end{lem}

\newcommand{\LS}{\mathit{LS}}
\newcommand{\GS}{\mathit{GS}}

\mypar{Sensitivity and Basic DP Mechanisms.} The main differentially private primitive we will use is \emph{smooth sensitivity}~\cite{NissimRS07}.  Let $f \from \cG_{n} \to \R$ be a real-valued function.  For a graph $G \in \cG_{n}$, we can define the \emph{local sensitivity} of $f$ at $G$ to be
$$
\LS_{f}(G) = \max_{G': G' \sim G} | f(G) - f(G') |
$$
and the \emph{global sensitivity} of $f$ to be
$$
\GS_{f} = \max_{G} \LS_{f}(G) = \max_{G' \sim G} | f(G) - f(G') |
$$

A basic result in differential privacy says that we can achieve privacy for any real-valued function $f$ by adding noise calibrated to the global sensitivity of $f$.
\begin{thm} [DP via Global Sensitivity~\cite{DworkMNS06}] \label{thm:gs}
Let $f : \cG_{n} \to \R$ be any function.  Then the algorithm
$$
\cA(G) = f(G) + \frac{\mathit{GS}_{f}}{\eps} \cdot Z,
$$
where $Z$ is sampled from a standard Laplace distribution, satisfies $(\eps,0)$-differential privacy.\footnote{The standard Laplace distribution $Z$ has $\ex{}{Z} = 0, \ex{}{Z^2} = 2$, and density $\mu(z) \propto e^{-|z|}$.}
Moreover, this mechanism satisfies $\ex{\cA}{(\cA(G) - f(G))^2} = O(\GS_{f}/\eps)$, and for all $t>0$ we have that
$$\pr{\cA}{|\cA(G) - f(G)| \geq t \cdot \GS_{f} / \eps} \leq \exp(-t).$$
\end{thm}

In many cases the global sensitivity of $f$ is too high, and we want to use a more refined mechanism that adds instance-dependent noise that is more comparable to the local sensitivity.  This can be achieved via the \emph{smooth sensitivity} framework of Nissim \etal~\cite{NissimRS07}.
\begin{definition}[Smooth Upper Bound~\cite{NissimRS07}]
Let  $f: \cG_{n} \to \R$ be a real-valued function and $\beta > 0$ be a parameter.
A function $S: \cG_{n} \to \R$ is a \emph{$\beta$-smooth upper bound on $\LS_f$}  if
\begin{enumerate}
\item for all $G \in \cG_{n}$, $S(G) \ge LS_f(G)$, and 
\item for all neighboring $G \sim G' \in \cG_{n}$, $S(G) \le e^\beta \cdot S(G')$. 
\end{enumerate}
\end{definition}

The key result in smooth sensitivity is that we can achieve differential privacy by adding noise to $f(G)$ proportional to any smooth upper bound $S(G)$.
\begin{thm}[DP via Smooth Sensitivity~\cite{NissimRS07,BunS19}] \label{thm:ss}
Let  $f: \cG_{n} \to \R$ be any function and $S$ be a $\beta$-smooth upper bound on the local sensitivity of $f$ for any $\beta \leq \eps$.  Then the algorithm 
$$\cA(G) = f(G) + \frac{S(G)}{\eps} \cdot Z,$$
where $Z$ is sampled from a Student's $t$-distribution with $3$ degrees of freedom, satisfies $(O(\eps),0)$-differential privacy.\footnote{The Student's $t$-distribution with 3 degrees of freedom can be efficiently sampled by choosing $X,Y_1,Y_2,Y_3 \sim \cN(0,1)$ independently from a standard normal and returning $Z = X / \sqrt{Y_1^2 + Y_2^2 + Y_3^2}$.  This distribution has $\ex{}{Z} = 0$ and $\ex{}{Z^2} = 3$, and its density is $\mu(z) \propto 1/(1+z^2)^2$.}
Moreover, for any $G \in \cG_{n}$, this algorithm satisfies $\ex{\cA}{(\cA(G) - f(G))^2} = O(S(G)^2/\eps^2)$.
\end{thm}

\section{An Estimator for Concentrated-Degree Graphs}

In this section we describe and analyze a node-differentially-private estimator for the edge density of a concentrated-degree graph.  

\subsection{The Estimator}

In order to describe the estimator we introduce some key notation.  The input to the estimator is a graph $G = (V,E)$ and a parameter $k^*$.  Intuitively, $k^*$ should be an upper bound on the concentration parameter of the graph, although we obtain more general results when $k^*$ is not an upper bound, in case the user does not have an \emph{a priori} upper bound on this quantity.

For a graph $G=(V,E)$, let $p_G = |E|/\binom{n}{2}$ be the empirical edge density of $G$, 
and let $\ad_G = (n-1)p_G$ be the empirical average degree of $G$.
Let $k_G$ be the smallest positive integer value such that at most $k_G$ vertices of $G$ have 
degrees differing from $\ad_G$ by more than $\kprime := \kstar + 3k_G$. 
Define $I_G = [\ad_G - \kprime, \ad_G + \kprime]$. 
For each vertex $v\in V$, let $t_v = \min\{|t| : \deg_G(v)\pm t \in I_G \}$ be the distance between $\deg_G(v)$ and the 
interval $I_G$, and define 
the \emph{weight} $\wt_G(v)$ of $v$ as follows. For a parameter $\beta > 0$ to be specified later, let
\[
\wt_G(v) = 
\begin{cases}
1 & \text{if } t_v = 0 \\
1 - \beta t_v & \text{if } t_v \in (0,1/\beta] \\
0 & \text{otherwise}.
\end{cases}
\]
That is, $\wt_G(v) = \max(0,1-\beta t_v)$.
For each pair of vertices $e=\{u,v\}$, define the \emph{weight} $\wt_G(e)$ and \emph{value} $\val_G(e)$ as follows. 
Let
\[
\wt_G(e) = \min(\wt_G(u), \wt_G(v))
\]
and let
\[
\val_G(e) = \wt_G(e) \cdot x_e + (1-\wt_G(e)) \cdot p_G
\]
where $x_e$ denotes the indicator variable on whether $e\in E$. 
As above, define the function $f$ to be the total value of all pairs of vertices in the graph,
\[
f(G) = 
\sum_{u,v \in V} \val_G(\{u,v\}),
\]
where the sum is over unordered pairs of distinct vertices.  

Once we construct this function $f$, we add noise to $f$ proportional to a $\beta$-smooth upper bound on the sensitivity of $f$, which we derive in this section.  Pseudocode for our estimator is given in Algorithm~\ref{algo:concentrated_degree}.

\adam{$\beta$ can be $\min(\epsilon, 1/\sqrt{k})$ or $\min(\epsilon, 1/k)$, it's essentially the same. 
(The first should be better for $k_G=0$, the second might become better as $k_G$ increases.}

\begin{algorithm}[t!] \label{algo:concentrated_degree} 
\caption{Estimating the edge density of a concentrated-degree graph.} \label{alg:cd}
\KwIn{A graph $G \in \cG_{n}$ and parameters $\eps > 0$ and $\kstar \ge 0$.} 
\KwOut{A parameter $0 \leq \hat{p} \leq 1$.} \vspace{10pt}

Let $p_G = \frac{1}{\binom{n}{2}} \sum_{e} x_{e}$ and $\ad_G = (n-1)p_G$.\\ \vspace{6pt}

Let $\beta = \min(\epsilon, 1/\sqrt{k})$.\\ \vspace{6pt}

Let $k_G > 0$ be the smallest positive integer such that at most $k_G$ vertices have degree outside 
$[\ad_G-\kstar - 3k_G, \ad_G + \kstar + 3k_G]$.\\ \vspace{6pt}

For $v\in V$, let $t_v = \min\{|t| : \deg_G(v)\pm t \in [\ad_G - \kstar - 3k_G, \ad_G + \kstar + 3k_G] \}$ and let
$\wt_G(v) = \max(0, 1-\beta t_v)$. \\ \vspace{6pt}

For each $u,v\in V$, let $\wt_G(\{u,v\}) = \min(\wt_G(u), \wt_G(v))$ and let $\val_G(e) = \wt_G(e) \cdot x_e + (1-\wt_G(e)) p_G$. \\ \vspace{6pt}

Let $\displaystyle f(G) = \sum_{u\neq v} \val_G(\{u,v\})$, where the sum is over unordered pairs of vertices.\\ \vspace{6pt}

Let $s = \max_{\ell \ge 0} c e^{-\beta \ell} \cdot  (k_G + \ell + \kstar + \beta (k_G+\ell)(k_G+\ell+\kstar) + 1/\beta)$,
where $c$ is the constant implied by Lemma~\ref{lem:LS_bound}.\\ \vspace{6pt}


Return $\frac{1}{\binom{n}{2}}\cdot \left(f(G) + (s/\epsilon)\cdot Z\right)$, where $Z$ is sampled from a Student's $t$-distribution with three degrees of freedom.\\

\end{algorithm}

\subsection{Analysis using Smooth Sensitivity}


We begin by bounding the local sensitivity $LS_f(G)$ of the function $f$ defined above. 

\begin{lemma}
\label{lem:LS_bound}
$LS_f(G) = O((k_G + \kstar)(1 + \beta k_G) + \frac{1}{\beta})$.
\end{lemma}
\begin{proof}
Consider any pair of graphs $G,G'$ differing in only a single vertex $v^*$, and
note that the empirical edge densities $p_G$ and $p_{G'}$ can differ by at most $\frac{2}{n} < \frac{2}{n-1}$, 
so $\ad_G$ and $\ad_{G'}$ can differ by at most $2$. 
Moreover, for any vertex $v\neq v^*$, the degree of $v$ can differ by at most $1$ between $G$ and $G'$. 
Consequently, by the Triangle Inequality, 
for any $v\neq v^*$, $|\ad_G - \deg_G(v)|$ can differ from $|\ad_{G'} - \deg_{G'}(v)|$ by at most $3$ 
and $\wt_G(v)$ can differ from $\wt_{G'}(v)$ by at most $3\beta$. 
It follows from the former statement that $k_G$ and $k_{G'}$ differ by at most $1$.

Let $\Far_G$ denote the set of at most $k_G$ vertices whose degree differs from $\ad_G$ by more than 
$\kprime = \kstar + 3k_G$. 
For any vertices $u,v \notin \Far_G \cup \Far_{G'}\cup \{v^*\}$, we have that $\wt_G(\{u,v\}) = \wt_{G'}(\{u,v\}) = 1$,
and so $\val_G(\{u,v\}) = \val_{G'}(\{u,v\})$, since the edge $\{u,v\}$ is present in $G$ if and only if it is present in $G'$. 

Now consider edges $\{u,v\}$ such that $u,v\neq v^*$ but $u\in \Far_G \cup \Far_{G'}$ (and $v$ may or may not be as 
well). 
If $\deg_G(u)\notin [\ad_G - \kprimeprime, \ad_G+\kprimeprime]$ for $\kprimeprime=\kprime + 1/\beta + 3$, then
$\wt_G(u) = \wt_{G'}(u) = 0$ and so $|\val_G(\{u,v\}) - \val_{G'}(\{u,v\})| = |p_G - p_{G'}| \le 2/n$.
Otherwise, $\deg_G(u) \in [\ad_G - \kprimeprime, \ad_G+\kprimeprime]$.
We can break up the sum
\[
f_u(G) := \sum_{v\neq u} \val_G(\{u,v\}) = \sum_{v\neq u} \wt_G(\{u,v\}) \cdot x_{\{u,v\}} + \sum_{v\neq u} (1-\wt_G(\{u,v\})) p_G.
\]
Since at most $k_G$ other vertices can have weight less than
the weight of $u$, we can bound the first term by
\[
\sum_{v\neq u} \wt_G(u) x_{\{u,v\}} \pm k_G \wt_G(u) 
= \deg_G(u) \wt_G(u) \pm k_G \wt_G(u)
\]
and the second term by
\[
p_G\cdot\left( (n-1) - \sum_{v\neq u} \wt_G(\{u,v\})\right)
= \ad_G - \ad_G \wt_G(u) \pm p_G k_G \wt_G(u).
\]
so the total sum is bounded by
\[
f_u(G) = \ad_G + (\deg_G(u) - \ad_G) \wt_G(u) \pm 2 k_G \wt_G(u).
\]
Since $|\wt_G(u) - \wt_{G'}(u)| \le 3\beta$, it follows that
\[
|f_u(G) - f_u(G')| \le 7 + 3\beta(\kprimeprime+3) + 9\beta + 6 \beta k_G 
= O(1 + \beta(k_G+\kstar)).
\]
\adam{Maybe expand this step}
Since there are at most $k_G+k_G' \le 2k_G+1$ vertices in $u\in \Far_G \cup \Far_{G'}\setminus \{v^*\}$,
the total difference in the terms of $f(G)$ and $f(G')$ corresponding to such vertices is at most 
$O(k_G + \beta k_G(k_G+\kstar)).$ However, we are double-counting any edges between two vertices in 
$u\in \Far_G \cup \Far_{G'}$; the number of such edges is $O(k_G^2)$, and for any such edge $e$,
$|\val_G(e) - \val_{G'}(e)| \pm O(\beta)$. Consequently the error induced by this double-counting is at most
$O(\beta k_G^2)$, so the total difference between the terms of $f(G)$ and $f(G')$ corresponding 
to such vertices is still $O(k_G + \beta k_G(k_G+\kstar)).$

Finally, consider the edges $\{u,v^*\}$ involving vertex $v^*$. If $\wt_G(v^*) = 0$ then 
\[ 
f_{v^*}(G) = \sum_{v\neq v^*} \val_G(\{v^*,v\}) = (n-1)p_G=\ad_G.
\]
If $\wt_G(v^*)=1$ then $\deg_G(v^*)\in [\ad_G-\kprime, \ad_G+\kprime]$,
so
\begin{align*}
f_{v^*}(G) &= \sum_{v\neq v^*} \val_G(\{v^*,v\})\\
&= \deg_G(v^*) \pm k_G \\
&= \ad_G \pm k' \pm k_G. \\
\end{align*}


Otherwise, $\deg_G(v^*)\in [\ad_G-\kprime - 1/\beta, \ad_G+\kprime + 1/\beta]$. 
Then we have that 
\begin{align*}
f_{v^*}(G) &= \sum_{v\neq v^*} \val_G(\{v^*,v\}) \\
&= \ad_G + (\deg_G(v^*)  - \ad_G) \wt_G(v^*) \pm k_G \wt_G(v^*)\\
&= \ad_G \pm (\deg_G(v^*) - \ad_G) \pm k_G,
\end{align*}
so in either case we have that $f_{v^*}(G) \in [\ad_G - O(k_G+\kstar+1/\beta), \ad_G + O(k_G+\kstar+1/\beta)]$.
Consequently
$|f_{v^*}(G) - f_{v^*}(G')| \le O(k_G + \kstar + 1/\beta)$. 

Putting everything together, we have that
$LS_f(G) = O((k_G + \kstar)(1+ \beta k_G)+ 1/\beta)$.
\end{proof}

We now compute a smooth upper bound on $LS_f(G)$. From the proof of Lemma~\ref{lem:LS_bound}, we have
that there exists some constant $C>0$ such that $LS_f(G) \le C((k_G + \kstar)(1 + \beta k_G) + \frac{1}{\beta})$.
Let $$g(k_G,\kstar,\beta)=C((k_G + \kstar)(1 + \beta k_G) + \tfrac{1}{\beta})$$ be this upper bound on $LS_f(G)$,
and
let
$$S(G) = \max_{\ell \ge 0} e^{-\ell \beta} g(k_G+\ell, \kstar, \beta).$$
\begin{lemma}
\label{lem:alg-ss}
$S(G)$ is a $\beta$-smooth upper bound on the local sensitivity of $f$. Moreover, 
$$S(G) = O((k_G + \kstar)(1 + \beta k_G)  + \tfrac{1}{\beta}).$$
\end{lemma}
\begin{proof}
For neighboring graphs $G,G'$,
we have that 
\begin{align*}
S(G') &= \max_{\ell \ge 0} e^{-\ell \beta} g(k_{G'}+\ell, \kstar, \beta)\\
&\le \max_{\ell \ge 0} e^{-\ell \beta} g(k_{G}+\ell+1, \kstar, \beta)\\
&= e^\beta \max_{\ell \ge 1} e^{-\ell \beta} g(k_{G}+\ell, \kstar, \beta)\\
&\le e^\beta \max_{\ell \ge 0} e^{-\ell \beta} g(k_{G}+\ell, \kstar, \beta)\\
&= e^\beta S(G).
\end{align*}
Moreover, for fixed $k_G, \kstar, \beta$, 
consider the function $h(\ell) = e^{-\ell \beta} g(k_G+\ell, \kstar \beta)$, and consider the derivative $h'(\ell)$.
We have that
\[
h'(\ell) = C \beta e^{-\ell \beta} (k_G + \ell)(1-\beta(k_G + \ell + \kstar)).
\]
Consequently the only possible local maximum for $\ell> 0$ would occur for $\ell = 1/\beta - k_G - \kstar$; note that
the function $h$ decreases as $\ell\to\infty$.
Consequently the maximum value of $h$ occurs for some $\ell \le 1/\beta$, and so
\begin{align*}
S(G) &= \max_{\ell \ge 0} h(\ell)\\
&= \max_{\ell \ge 0} ce^{-\ell \beta} (k_G + \ell + \kstar +  (k_G+\ell)(k_G+\ell+\kstar)\beta + 1/\beta)\\
&\le C \cdot(k_G + 1/\beta + \kstar +  (k_G+1/\beta)(k_G+1/\beta+\kstar)\beta + 1/\beta)\\
&= C \cdot (3k_G  + 2\kstar + \beta k_G (k_G + \kstar) + 3/\beta)\\
&= O((k_G + \kstar)(1 + \beta k_G) + 1/\beta)
\end{align*}
as desired.
\end{proof}

\begin{theorem}
Algorithm~\ref{alg:cd} is $(O(\epsilon),0)$-differentially private. Moreover, for any $k$-concentrated $n$-vertex 
graph $G=(V,E)\in\cG_{n,k}$ with $k\ge 1$, we have that Algorithm~\ref{alg:cd} satisfies
\[
\ex{\cA}{\left(\frac{|E|}{\binom{n}{2}} - \cA_{\epsilon,k}(G)\right)^2\;} = O\left( \frac{k^2}{\epsilon^2 n^4} + \frac{1}{\epsilon^4 n^4}   \right)
\]
\end{theorem}
\begin{proof}
Algorithm~\ref{alg:cd} computes function $f$ and releases it with noise proportional to a $\beta$-smooth 
upper bound on the local sensitivity for $\beta \le \epsilon$. Consequently  $(O(\epsilon),0)$-differential privacy
follows immediately from Theorem~\ref{thm:ss}. 

We now analyze its accuracy on $k$-concentrated graphs $G$.
If $G$ is $k$-concentrated and $k^* \ge k$, then $\wt_G(v)=1$ for all vertices $v\in V$ and $\val_G(\{u,v\}) = x_{\{u,v\}}$ for all
$u,v\in V$, and so $f(G) = |E|$. Consequently Algorithm~\ref{alg:cd} computes the edge density
of a $k$-concentrated graph with noise distributed according to the Student's $t$-distribution scaled by a factor of
$S(G) / (\epsilon \binom{n}{2})$. 

Since $G$ is $k$-concentrated, we also have that $k_G = 1$, and so 
$S(G) = O(k + \beta(k+1) + 1/\beta) \le O(k + 1/\epsilon)$ by Lemma~\ref{lem:alg-ss}.
The variance of the Student's $t$-distribution with three degrees of freedom is $O(1)$,
so the expected squared error of the algorithm is
\[
O\left( \frac{(k+1/\epsilon)^2}{\epsilon^2 n^4}\right)
= O\left( \frac{k^2}{\epsilon^2  n^2} + \frac{1}{\epsilon^4 n^4}\right)
\] 
as desired.
\end{proof}

\section{Application to \ER~Graphs}

In this section we show how to apply Algorithm~\ref{alg:cd} to estimate the parameter of an \ER~graph.
Pseudocode is given in Algorithm~\ref{alg:er}.

\begin{algorithm} \label{algo:put_ineff}
\caption{Estimating the parameter of an \ER~graph.} \label{alg:er}
\KwIn{A graph $G \in \cG_{n}$ and parameters $\eps,\alpha > 0$.}
\KwOut{A parameter $0 \leq \hat{p} \leq 1$.} \vspace{10pt}

Let $\tilde{p}'
 \gets \frac{1}{\binom{n}{2}} \sum_{e} x_{e} + (2/\eps n) \cdot Z$ where $Z$ is a standard Laplace \\ \vspace{6pt}
Let $\tilde{p} \gets \tilde{p}' + 4 \log(1/\alpha) / \eps n$ and $\tilde{k} \gets \sqrt{\tilde{p} n \log (n/\alpha)}$ \\ \vspace{8pt}
Return $\hat{p} \gets \cA_{\tilde{k},\eps}(G)$ where $\cA_{\tilde{k},\eps}$ is Algorithm~\ref{alg:cd} with parameters $\tilde{k}$ and $\eps$ \\
\end{algorithm}

It is straightforward to prove that this mechanism satisfies differential privacy.
\begin{thm}
Algorithm~\ref{alg:er} satisfies $(O(\eps),0)$-node-differential privacy.
\end{thm}
\begin{proof}
The first line computes the empirical edge density of the graph $G$, which is a function with global sensitivity $(n-1)/\binom{n}{2} = 2/n$.  Therefore by Theorem~\ref{thm:gs} this step satisfies $(\eps,0)$-differential privacy.  The third line runs an algorithm that satisfies $(O(\eps),0)$-differential privacy for every fixed parameter $\tilde{k}$.  By Lemma~\ref{lem:dpcomp}, the composition satisfies $(O(\eps),0)$-differential privacy.
\end{proof}

Next, we argue that this algorithm satisfies the desired accuracy guarantee.
\begin{thm}
For every $n \in \N$ and $\frac12 \geq p \geq 0$, and an appropriate parameter $\alpha > 0$, Algorithm~\ref{alg:er} satisfies
$$
\ex{G \sim G(n,p), \cA}{(p - \cA(G))^2} = \frac{p(1-p)}{\binom{n}{2}} + \tilde{O}\left( \frac{\max\{p,\frac1n\}}{\eps^2 n^3} + \frac{1}{\eps^4 n^4}   \right)
$$
\end{thm}
\begin{proof}
We will prove the result in the case where $p \geq \frac{\log n}{n}$.  The case where $p$ is smaller will follow immediately by using $\frac{\log n}{n}$ as an upper bound on $p$.  The first term in the bound is simply the variance of the empirical edge-density $\bar{p}$.  For the remainder of the proof we will focus on bounding $\ex{}{(\bar{p} - \hat{p})^2}$.

A basic fact about $G(n,p)$ for $p \geq \frac{\log n}{n}$ is that with probability at least $1-2\alpha$: (1) $|\bar{p} - p| \leq 2\log(1/\alpha)/n$, and
(2) the degree of every node $i$ lies in the interval $[ \ad \pm \sqrt{pn \log(n/\alpha)} ]$ where $\ad$ is the average degree of $G$.  
We will assume for the remainder that these events hold.

Using Theorem~\ref{thm:gs}, we also have that with probability at least $1-\alpha$, the estimate $\tilde{p}'$ satisfies $| \bar{p} - \tilde{p}' | \leq 4 \log(1/\alpha) / \eps n$.  We will also assume for the remainder that this latter event holds.  Therefore, we have
$p \leq \tilde{p}$ and $p \geq \tilde{p} - 8 \log(1/\alpha)/\eps n.$

Assuming this condition holds, the graph will have $\tilde{k}$-concentrated degrees for $\tilde{k}$ as specified on line 2 of the algorithm.  Since this assumption holds, we have
\begin{align*}
\ex{}{(\bar{p} - \cA_{\tilde{k},\eps}(G))^2} 
={} &\tilde{O}\left( \frac{\tilde{k}^2}{\eps^2 n^4} + \frac{1}{\eps^4 n^4} \right)\\ 
={} &\tilde{O}\left( \frac{\tilde{p} n}{\eps^2 n^4} + \frac{1}{\eps^4 n^4} \right) \\
={} &\tilde{O}\left( \frac{p n + \frac{1}{\eps n}}{\eps^2 n^4} + \frac{1}{\eps^4 n^4} \right)\\
={} &\tilde{O}\left( \frac{p n}{\eps^2 n^4} + \frac{1}{\eps^4 n^4} \right).
\end{align*}

To complete the proof, we can plug in a suitably small $\alpha = 1/\mathrm{poly}(n)$ so that the $O(\alpha)$ probability of failure will not affect the overall mean-squared error in a significant way.
\end{proof}

\section{Lower Bounds for Concentrated-Degree Graphs}

In this section we prove a lower bound for estimating the number of edges in concentrated-degree graphs.  Theorem~\ref{thm:main-lb}, which lower bounds the mean squared error follows by applying Jensen's Inequality.
\begin{thm} \label{thm:main-lb-full}
For every $n,k \in \N$, every $ \eps \in [\frac{2}{n},\frac14]$ and $\delta \leq \frac{\eps}{32}$, and every $(\eps,\delta)$-node-DP algorithm $A$, there exists $G \in \cG_{n,k}$ such that
$
\ex{A}{ | p_{G} - A(G) |} = \Omega\left( \frac{k}{\eps n^2} + \frac{1}{\eps^2 n^2} \right).
$ 
\end{thm}

The proof relies on the following standard fact about differentially private algorithms.
Since we are not aware of a formal treatment in the literature, 
we include a proof for completeness.
\begin{lem} \label{lem:locallb}
Suppose there are two graphs $G_0, G_1 \in \cG_{n,k}$ at node distance at most $\frac{1}{\eps}$ from one another.  Then for every $(\eps,\frac{\eps}{32})$-node-DP algorithm $A$, there exists $b \in \{0,1\}$ such that
$$
\ex{A}{\left| p_{G_{b}} - A(G_{b}) \right|} = \Omega\left(  \left| p_{G_{0}} - p_{G_{1}} \right|  \right).
$$
\end{lem}
\begin{proof}
Let $A$ be any $\eps$-node-DP algorithm.  Since $G_0,G_1$ have node distance at most $\frac{1}{\eps}$, by group privacy (Lemma~\ref{lem:groupdp}), for every set $S$ and every $b \in \{0,1\}$
$$
\pr{}{A(G_b) \in S} \leq e \cdot \pr{}{A(G_{1-b}) \in S} + \tfrac{1}{16}.
$$
Now, let $S_{b} = \set{y : \left| y - p_{G_{b}} \right| < \frac12 \left| p_{G_{0}} - p_{G_{1}} \right|}$ and note that $S_{0}$ and $S_{1}$ are disjoint by construction.  Let $\rho = \min\{ \pr{}{A(G_{0}) \in S_{0}}, \pr{}{A(G_1) \in S_{1}}\}$.  Then we have
\begin{align*}
1-\rho \geq{} &\pr{}{A(G_0) \not\in S_0} \\
\geq{} &\pr{}{A(G_0) \in S_1} \\
\geq{} &e^{-1} \pr{}{A(G_1) \in S_1} - \tfrac{1}{16} \\
\geq{} &e^{-1} \rho - \tfrac{1}{16}
\end{align*}
from which we can deduce $\rho \leq \frac{4}{5}$.  Therefore, for some $b \in \zo$, we have 
$$
\pr{}{ | p_{G_{b}} - A(G_{b}) | \geq \tfrac12 \left| p_{G_{0}} - p_{G_{1}} \right|  } 
\geq \tfrac{1}{5},
$$
from which the lemma follows.
 \end{proof}

We will construct two simple pairs of graphs to which we can apply Lemma~\ref{lem:locallb}.
\begin{lem}[Lower bound for large $k$] \label{lem:largek}
For every $n, k \in \N$ and $\eps \geq 2/n$, there is a pair of graphs $G_0,G_1 \in \cG_{n,k}$ at node distance $1/\eps$ such that $| p_{G_{0}} - p_{G_{1}} | = \Omega(\frac{k}{\eps n^2})$.
\end{lem}
\begin{proof}
Let $G_0$ be the empty graph on $n$ nodes.  Note that $p_{G_0} = 0$, $\ad_{G_0} = 0$, and $G_0$ is in $\cG_{n,k}$. 

We construct $G_1$ as follows.  Start with the empty bipartite graph with $\frac{1}{\eps}$ nodes on the left and $n - \frac{1}{\eps}$ nodes on the right.  We connect the first node on the left to each of the first $k$ nodes on the right, then the second node on the left to each of the next $k$ nodes on the right and so on, wrapping around to the first node on the right when we run out of nodes.  By construction, $p_{G_1} = k / \eps \binom{n}{2}$, $\ad_{G_{1}} =  2k / \eps n$.  Moreover, each of the first $\frac{1}{\eps}$ nodes has degree exactly $k$ and each of the nodes on the right has degree
$$
\frac{k / \eps}{n - 1/\eps} \pm 1 = \frac{k}{\eps n - 1} \pm 1
$$
Thus, for $n$ larger than some absolute constant, every degree lies in the interval $[\ad_{G_{1}} \pm k]$ so we have $G_{1} \in \cG_{n,k}$.  \end{proof}

\begin{lem}[Lower bound for small $k$] \label{lem:smallk}
For every $n\ge 4$ and $\eps\in[2/n,1/4]$,
there is a pair of graphs $G_0,G_1 \in \cG_{n,1}$ at node distance $1/\eps$ such that $| p_{G_{0}} - p_{G_{1}} | = \Omega(\frac{1}{\eps^2 n^2})$.
\end{lem}
\begin{proof}
Let $i = \lceil n \epsilon \rceil$, and let $G_0$ be the graph consisting of $i$ disjoint cliques each of size 
$\lfloor n/i \rfloor$ or $\lceil n/i \rceil$. Let $G_1$ be the graph consisting of $i+1$ disjoint cliques each of size 
$\lfloor n/(i+1) \rfloor$ or $\lceil n/(i+1) \rceil$. We can obtain $G_0$ from $G_1$ by taking one of the cliques and
redistributing its vertices among the $i$ remaining cliques, so $G_0$ and $G_1$ have node distance
$\ell := \lfloor n/(i+1) \rfloor \le 1/\epsilon$. 
For $1/4 \ge \epsilon \ge 2/n$ we have that $\ell \ge \lfloor 1/2\epsilon \rfloor > 1/4\epsilon$.
Transforming $G_1$ into $G_0$ involves removing a clique of size $\ell$, containing $\binom{\ell}{2}$ edges, 
and then inserting these $\ell$ vertices into cliques that already have size $\ell$, adding at least $\ell^2$ new edges.
Consequently $G_0$ contains at least $\ell^2 - \ell(\ell-1)/2 = \ell(\ell+1)/2$ more edges than $G_1$,
so 
\[
|p_{G_1} - p_{G_0}| \ge \frac{\binom{\ell+1}{2}}{\binom{n}{2}}
\ge \frac{\ell^2}{n^2} \ge \Omega(1/\eps^2 n^2),
\] 
as desired.
\end{proof}

Theorem~\ref{thm:main-lb-full} now follows by combining Lemmas~\ref{lem:locallb},~\ref{lem:largek}, and~\ref{lem:smallk}.

\section*{Acknowledgments}

\noindent Part of this work was done while the authors were visiting the Simons Institute for the Theory of Computing.
AS is supported by NSF MACS CNS-1413920, DARPA/NJIT Palisade 491512803, Sloan/NJIT 996698, and MIT/IBM W1771646.  JU is supported by NSF grants CCF-1718088, CCF-1750640, and CNS-1816028.  The authors are grateful to Adam Smith for helpful discussions.  

\bibliographystyle{alpha}
\bibliography{./bibliography/refs}

\end{document}